The reaction of liquid silicon tetrachloride and magnesium forming silicon and  magnesium chloride. This is one step in the preparation of ultrapure silicon used  in the semiconductor industry.


1
Expert's answer
2021-11-29T04:37:48-0500

(a)

SiCl4 (s) + 2 Mg (s) -----------> Si (s) + 2 MgCl2 (s)

(b)

Mass of Si = 0.130 kg = 130 g.

Molar mass of Si = 28.1 g/mol

Moles of Si = mass / molar mass = 130 / 28.1 = 4.63 mol

From the balanced equation,

1 mol o fSi is obtained from 1 mol of SiCl4

Then, 4.63 mol of Si is obtained from 4.63 mol of SiCl4

Mass of SiCl4 = moles * molar mass = 4.63 * 170.1 = 788 g. = 0.788 kg. of SiCl4 is need.

From the balanced equation,

1 mol of Si is obtained from 2 mol of Mg

Then,

4.63 mol of Si is obtained from 4.63 * 2 = 9.26 mol of Mg

Mass of Mg = 9.26 * 24 = 222. g. = 0.222 kg. of Mg is neded.

(c)

Since actual yield is 93.0 %,

Mass of SiCl4 = 0.788 * 100 / 93 = 0.847 kg.

Mass of Mg = 0.222 * 100 / 93 = 0.239 kg.